# PhyGDPO: Physics-Aware Groupwise Direct Preference Optimization for Physically Consistent Text-to-Video Generation

### v2

@@ 3 Experiment
Table 2: Ablation study on VideoPhy2 [4]. (a) and (b) study the components of PhyAugPipe and PhyGDPO towards better performance. (c) compares VideoDPO with PhyGDPO with the same settings. (d) studies the effect of LoRA-SR scheme.
Figure 7: Visual analysis. (a) When using LoRA-SR, groupwise probabilistic modeling, and physics-guided rewarding, the generated tennis ball motion becomes more physically realistic, better adhering to the laws of buoyancy. (b) Using our PhyGDPO can generate more physically coherent actions with well-structured human body motion than VideoDPO and Flow-DPO. (c) Applying our LoRA-SR can generate more accurate and plausible hand–shovel contact in snow shoveling.
@@ 3.1 Comparison with State-of-the-Art Methods
User Study. Since the auto-evaluation tools of VideoPhy2 and PhyGenBench are based on VLMs, their assessment of physical plausibility may be inaccurate. Therefore, we conduct a user study with 104 participants. Each participant views the videos generated by our PhyGDPO and a random competing method, along with the randomly selected text prompt from VideoPhy2 and PhyGenBench. The participants are asked to pick up the video that better follows the physics laws. Tab. LABEL:tab:user_study reports the user preference (%) of our method over competing entries. PhyGDPO is consistently preferred by the human evaluators than all SOTA methods.
@@ 4.1 Text-to-Video Generation Models
T2V generation [6, 68, 24, 74, 5, 23, 5, 17, 27, 25, 71, 59] has witnessed significant progress. Many existing works follow DiT [46] to adopt a Transformer [53] to predict the noise in diffusion [41, 34, 39, 40, 20, 48, 8, 22, 32] or estimate the velocity field in flow matching [55, 27, 17, 21, 7, 28, 72, 56, 70, 31, 13]. Although high visual quality is achieved, accurately modeling the underlying physics-related effects remains challenging. Two recent works, DiffPhy [69] and PhyT2V [64], adopt an LLM to extend the text prompt with explicit physics laws and phenomena, and then iteratively generates the video or finetunes a T2V model with extended prompts. Yet, these prompt extension-based methods are easily misled by the mistakes of LLM and struggle to learn the implicit physics. Different from these works, we aim to explore the physics reasoning potential of the T2V model itself.
@@ 4.2 Direct Preference Optimization
Direct Preference Optimization (DPO) [49] is proposed to align LLMs with human preferences and has been adopted in image [54, 43, 16, 30] and video [36, 12, 35, 63, 67] generation. For example, DiffusionDPO [54] finetunes a text-to-image diffusion model with human-annotated preference image pairs. VideoDPO [36] adapts DiffusionDPO into T2V generation. However, these DPO frameworks mainly improve the visual quality and aesthetics. In addition, these DPO methods suffer from a severe low-efficiency problem because they need to copy a full model as the reference and they also lack sufficient guidance of physics, which makes the preference signal less reliable and limits the improvement on physical consistency. Our work aims to cope with these research problems.

### v3

@@ 3 Experiment
Evaluation Metrics. VideoPhy2 uses another VLM - VideoPhy2-AutoRater [5], which is different from VideoCon-Physics [4], to evaluate the semantics adherence and physics commonsense of the videos with scores ranging from 1 to 5, and then compute the ratio with both scores $\geq$ 4. PhyGenBench evaluation is based on GPT-4o [1], CLIP [52], InternVideo2 [63], and VideoLLaVA [37]. It firstly detects key physical phenomena required by the prompt and then assesses their order and naturalness, covering 27 physical laws in 4 basic physics domains.
@@
Method PhyT2V VideoDPO Wan2.1-14B Veo3.1 Sora2 PhyGDPO
Overall 0.3186 0.3288 0.4119 0.5034 0.5373 0.5525
Method Vcrafter2 + Flow-DPO + VideoDPO + PhyGDPO
Overall 0.1034 0.1128 0.1373 0.1426
Table 2: Ablation study on VideoPhy2 [5]. (a) and (b) study the components of PhyAugPipe and PhyGDPO towards better performance. (c) compares VideoDPO with PhyGDPO with the same settings. (d) studies the effect of LoRA-SR scheme. (e) and (f) are cross evaluation by changing the VLM scorer to Gemini-2.5-pro [17] and T2V base model to Vcrafter2 [12].
Figure 7: Break-down visual analysis of PhyGDPO. When we progressively apply LoRA-SR, groupwise DPO modeling, and physics rewarding to Wan2.1-14B, the tennis ball in the generated video increasingly conforms to the physical law of buoyancy, floating stably on the water surface instead of sinking down or bouncing unrealistically above it.
@@ 3.1 Comparison with State-of-the-Art Methods
User Study. Since the auto-evaluation tools of VideoPhy2 and PhyGenBench are based on VLMs, their assessment of physical plausibility may not be perfect. Thus, we conduct a user study with 104 participants. Each participant is asked to pick up the video that better follows the physical laws. Tab. LABEL:tab:user_study reports the user preference (%) of our method over competing entries. Each participant completes 48 trials (one prompt per trial), with 24 prompts randomly sampled from VideoPhy2 and 24 from PhyGenBench. Each trial randomly selects one of the 8 baselines (non-repeated) in Tab. LABEL:tab:user_study and compares it to our method given the same prompt. Video results are presented anonymously in random order. All results in Tab. LABEL:tab:user_study have 95% confidence intervals no wider than $\pm 2.4\%$. PhyGDPO is preferred by human evaluators, indicating that PhyGDPO learns genuine physics beyond VLM agreement.
$\alpha_{\text{min}}$ 0.3 0.4 0.5 0.6 0.7
Overall 0.1582 0.1591 0.1627 0.1615 0.1608
$k_{\alpha}$ 4.0 4.5 5.0 5.5 6.0
Overall 0.1611 0.1579 0.1627 0.1593 0.1618
$b_{\alpha}$ 0.3 0.4 0.5 0.6 0.7
Overall 0.1582 0.1579 0.1627 0.1598 0.1602
$\lambda$ 0.3 0.4 0.5 0.6 0.7
Overall 0.1592 0.1586 0.1608 0.1627 0.1615
$k_{\gamma}$ 1.0 1.5 2.0 2.5 3.0
Overall 0.1588 0.1609 0.1627 0.1592 0.1598
$b_{\gamma}$ 0.3 0.4 0.5 0.6 0.7
Overall 0.1586 0.1627 0.1582 0.1607 0.1595
Table 3: Analysis of the hyperparameters in the physics rewarding of Eq.(9) on the VideoPhy2 [5] benchmark. When tweaking a hyperparameter to analyze its effectiveness, other hyperparameters are fixed at their optimal values.
Figure 8: Visual comparison with state-of-the-art DPO algorithms on Wan2.1-T2V-1.3B under the same training settings and data for fairness. Using our PhyGDPO can generate more physically plausible video with undistorted human body motion.
@@ 3.2 Ablation Study
Figure 9: Visual analysis of our LoRA-SR scheme. Applying our LoRA-SR can generate more accurate and plausible hand–shovel interaction in the snow shoveling scenario. In contrast, removing LoRA-SR generates distorted hand actions.
@@
Cross Evaluation. (i) To provide stronger evidence that our method improves the physical plausibility, we conduct a cross-VLM evaluation. We replace the VLM scorer of VideoPhy2 by a stronger physics evaluator, Gemini-2.5-pro [17], to score the generated videos with the same protocol in Tab. LABEL:tab:cross-vlm. Our method still outperforms SOTA T2V models, showing that our method learns better genuine physics. (ii) We change the T2V model to Vcrafter2 [12] and apply different DPO methods on it with the same setting for fairness in Tab. LABEL:tab:cross_t2v. PhyGDPO consistently surpasses SOTA DPO methods, suggesting the superiority and generalization ability of our method.
Parameter Analysis. We analyze the effects of the hyperparameters in Eq. (9). When tweaking a hyperparameter, other hyperparameters are fixed at their optimal values. The results are shown in Tab. 3. The best performance is achieved when $\alpha_{\text{min}}=0.5$, $k_{\alpha}=5.0$, $b_{\alpha}=0.5$, $\lambda=0.6$, $k_{\gamma}=2.0$, and $b_{\gamma}=0.4$. As shown in Tab. LABEL:tab:phyaugpipe, when we do not use the physics-guided rewarding, the model achieves 0.1575 on the overall score. And across all hyperparameter configurations of Tab. 3, our method consistently achieves performance gains.
@@ 4.1 Text-to-Video Generation Models
T2V generation [7, 72, 26, 78, 6, 25, 6, 19, 29, 27, 75, 62] has witnessed significant progress. Many existing works follow DiT [49] to adopt a Transformer [56] to predict the noise in diffusion [44, 36, 42, 43, 22, 51, 9, 24, 34] or estimate the velocity field in flow matching [58, 29, 19, 23, 8, 30, 76, 59, 74, 33, 14]. Although high visual quality is achieved, accurately modeling the underlying physics-related effects remains challenging. Two recent works, DiffPhy [73] and PhyT2V [68], adopt an LLM to extend the text prompt with explicit physics laws and phenomena, and then iteratively generates the video or finetunes a T2V model with extended prompts. Yet, these prompt extension-based methods are easily misled by the mistakes of LLM and struggle to learn implicit physics. We aim to fill this gap.
@@ 4.2 Direct Preference Optimization
Direct Preference Optimization (DPO) [53] is proposed to align LLMs with human preferences and has been adopted in image [57, 46, 18, 32] and video [39, 13, 38, 67, 71] generation. For example, DiffusionDPO [57] finetunes a text-to-image diffusion model with human-annotated preference image pairs. VideoDPO [39] adapts DiffusionDPO into T2V generation. However, these DPO frameworks mainly improve the visual quality and aesthetics. In addition, these DPO methods suffer from a severe low-efficiency problem because they need to copy a full model as the reference, which occupies redundant GPU memory. Our work focused on solving these research problems.